At the beginning of the year, Oakmont Company bought three used machines from American Manufacturing, Inc. The machines immedRequired 1 Required 2 Compute the cost of each machine Cost of Machine Machine A Machine B Machine CJournal entry worksheet Record the depreciation expense for the three used machines at the end of year 1 assuming each machin
Answer #1

Requirement 1

Total Cost of Machine
Machine A $    22,200.00
Machine B $    13,000.00
Machine C $    11,900.00

Working

Amount paid for assets Installation cost Renovation cost Total Cost of Machine
Machine A $       20,700.00 $            850.00 $               650.00 $    22,200.00
Machine B $       11,200.00 $        1,000.00 $               800.00 $    13,000.00
Machine C $       10,900.00 $            300.00 $               700.00 $    11,900.00

Requirement 2

Transaction General Journal Debit Credit
1 Depreciation expense $       8,780.00
           Accumulated depreciation-Machine A $        4,120.00
           Accumulated depreciation-Machine B $        2,280.00
           Accumulated depreciation-Machine C $        2,380.00
(To record depreciation expense)

.

Working

Straight line Method
A Cost $ 22,200
B Residual Value $ 1,600
C=A - B Depreciable base $ 20,600
D Life [in years left ]                                  5
E=C/D Annual SLM depreciation $               4,120.00

.

Units of Production method
A Cost $ 13,000
B Residual Value $ 1,600
C=A - B Depreciable base $ 11,400
D Usage in units(in Hours) 20000
E Depreciation per hour $                        0.57

.

Double declining Method
A Cost $ 11,900
B Residual Value $ 1,600
C=A - B Depreciable base $ 10,300
D Life [in years] 10
E=C/D Annual SLM depreciation $ 1,030
F=E/C SLM Rate 10.00%
G=F x 2 DDB Rate 20.00%

.

Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
1 $ 11,900 20.00% $ 2,380 $ 2,380 $ 9,520